OpenStudy (anonymous):

5/8 + 2/7 if possible simplify please show all work , write final answer ????

OpenStudy (anonymous):

Do you know What is LCM or LCD and how we take LCD of two numbers..??

OpenStudy (anonymous):

i need help

OpenStudy (anonymous):

find common denominator which would be 56 so then time 5/8 by 7 and 2/7 by 8 5/8 *7=35/56 2/7*8=16/56 now you add them 51/56

OpenStudy (anonymous):

51/56

jimthompson5910 (jim_thompson5910):

candygirl7655 and MissSosa are correct

OpenStudy (anonymous):

can someoe help me?

OpenStudy (anonymous):

Okay you guys one more question and you four are life savers

OpenStudy (anonymous):

whats the question